OpenStudy (anonymous):

You are creating an emergency fund and decide to place $415/month in an account that earns a 3.75%APR. How much interest accrues by the end of the first full month? I don't know if I did it right, but the answer I got was..... $129.69 Anyone know if that is right?

OpenStudy (anonymous):

I don't think I did it right!

OpenStudy (unklerhaukus):

what does APR stand for?

OpenStudy (anonymous):

Annual Percentage Rate? I think.

jimthompson5910 (jim_thompson5910):

They want the interest, not the full amount in the account

OpenStudy (anonymous):

I understand that. I guess I just don't understand how to get the interest amount. Can you explain the process of it? You don't have to give the answer, but can you explain the steps?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

Use the formula I = Prt

jimthompson5910 (jim_thompson5910):

P = 415 r = 0.0375 t = 1/12 (since t is in years)

OpenStudy (anonymous):

Okay, so it would be.... $1.30?

OpenStudy (anonymous):

Yep.

OpenStudy (anonymous):

You had to divide the 3.75% p.a by 12, to find the interest rate per month.

OpenStudy (anonymous):

Okay, I have another one to do, so i'm going to do it real quick, and can you check it for me?! (:

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

Got to go mate. This was the last answer I was going to give. @jim_thompson5910 may continue to help you if he's not busy. Have fun mate.

OpenStudy (anonymous):

You consistently deposit $250.00 into a savings account on the 15th of each month, and the account earns a 2.5% APR. How much is the balance of your savings account at the end of the 3rd month? Okay, so I got the interest total to be.. $1.56 Now do I add that to the $250.00?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

Wait, I would do 250*3 to get all three months right?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

I got 751.56 as the answer.

jimthompson5910 (jim_thompson5910):

this sounds like an annuity problem

jimthompson5910 (jim_thompson5910):

have you heard of that term?

OpenStudy (anonymous):

Yes, I have.

jimthompson5910 (jim_thompson5910):

just trying to figure out if it's an ordinary annuity or an annuity due

OpenStudy (anonymous):

No idea.

jimthompson5910 (jim_thompson5910):

well let's assume it's an ordinary annuity

jimthompson5910 (jim_thompson5910):

So you would use this formula FV = C [ ( (1+i)^n - 1)/i ] where in this case C = 250 i = 0.025/12 = 0.00208 n = 3

OpenStudy (anonymous):

I don't even know how to begin to use that formula.

jimthompson5910 (jim_thompson5910):

just plug in the given values and use a calculator to evaluate